# Brno, Czech Republic

### v1

@@ Abstract
We introduce the AnnoPage Dataset, a novel collection of 7 55075507\,5507 550 pages from historical documents, primarily in Czech and German, spanning from 1485 to the present, focusing on the late 19th and early 20th centuries. The dataset is designed to support research in document layout analysis and object detection. Each page is annotated with axis-aligned bounding boxes (AABB) representing elements of 25 categories of non-textual elements, such as images, maps, decorative elements, or charts, following the Czech Methodology of image document processing. The annotations were created by expert librarians to ensure accuracy and consistency. The dataset also incorporates pages from multiple, mainly historical, document datasets to enhance variability and maintain continuity. The dataset is divided into development and test subsets, with the test set carefully selected to maintain the category distribution. We provide baseline results using YOLO and DETR object detectors, offering a reference point for future research. The AnnoPage Dataset is publicly available on Zenodo [^1] https://doi.org/10.5281/zenodo.12788419, along with ground-truth annotations in YOLO format.
@@
Figure 1: Example pages from the AnnoPage Dataset.
@@ 1 Introduction
In recent years, the digitization of documents in various institutions has made the content of documents accessible to the general public. However, while the textual content of documents is usually easily recognizable using today’s text recognition technologies (OCR), other document content is often not processed, even though it may contain important information. Such content includes images, photographs, charts, various diagrams and drawings, musical notation, etc. The first and key step in processing such elements is the ability to detect them on the page and determine their category. Subsequently, these elements can be processed according to their category; for instance, images can be indexed for search, musical notations interpreted using optical music recognition (OMR), and mathematical expressions analyzed through math recognition systems. Such category-specific processing greatly expands the possibilities for users interacting with digitised documents, benefiting both general audiences and specialists conducting targeted research.
For document layout analysis and object detection in documents, many datasets exist [22, 31, 1, 28, 11, 8, 7, 17, 25]. However, these datasets contain only several categories of elements [31, 20, 7], they focus mainly on different parts of the text, such as captions, titles, headers, footers, etc. [22, 18, 17], they are synthetically generated [20], or they focus only on modern documents [1, 31, 22, 18, 17]. Some datasets focus only on a single category, such as text lines [8, 14] or tables [24, 25, 16, 10]. Also, historical documents, preserved by numerous institutions, also contain unique visual elements typically absent from modern documents, such as friezes, vignettes, signets, and other decorative components.
In this paper, we present the novel AnnoPage Dataset, which contains 7 55075507\,5507 550 pages from various, mainly historical documents. The pages were written mostly in Czech and German at the end of the 19th and during the 20th century (see Figure 1 for example pages). These pages were carefully annotated by librarians who used axis-aligned bounding boxes to identify elements of 25 categories of non-textual elements according to the Czech Methodology of image document processing [12].
@@ 2 Related work
In general, document layout analysis is a task for which a large number of diverse datasets exist [14, 8, 26, 11, 22, 31]. There are also datasets that focus on detecting various non-textual elements such as images, tables, or mathematical formulas [1, 7, 3, 6, 20, 4, 28, 31, 22, 18, 17]. In the terms of annotations, the ground-truth typically contains elements annotated either as bounding boxes [3, 4, 28, 31, 22, 16, 18] or bounding polygons [1, 7, 20, 6]. One of the existing datasets designed for document layout analysis that has a different type of annotation is the HBA 1.0 dataset [19], which contains pixel-level annotations.
@@ 3.1 Data acquisition
We have collected 7 55075507\,5507 550 pages from various documents. Most of the pages (5 690) come from publicly accessible documents from the Czech Digital Library [^2] https://www.digitalniknihovna.cz/. These pages are written mostly in Czech and German, but there are also a few pages containing other languages, such as French, and English. In terms of dating, the pages are from as early as 1485 to the present, with the majority from the late 19th and the first half of the 20th centuries. The distribution of the pages over the years is shown in Figure 2.
The rest of the pages come from existing datasets, most of them containing historical pages, namely IlluHisDoc [20] (237 pages), PRImA Layout Analysis Dataset [1] (170 pages), PRImA RDCL2019 [6] (40 pages), PRImA Europeana Newspapers [7] (361 pages), ICDAR2019 cBAD dataset [8] (382 pages), and TexBiG [28] (670 pages). The purpose of these extra pages is to increase the variability of the final dataset.
@@ 3.2 Element categories
Figure 3: Examples of selected element categories, which are less common and mostly related to historical documents.
@@ 3.4 Final dataset
The final dataset contains 27 9042790427\,90427 904 annotated elements on 6 72667266\,7266 726 pages and 824 pages without any annotated element. Each element is represented by a tightly fitted axis-aligned bounding box (AABB). The distribution of the number of annotations for each category is shown in Figure 4(a). The dataset is divided into development and test subsets. The further division of the development subset into the traditional training and validation subsets is left to the users. The test subset contains 600 pages (21 pages without any annotated element) coming from the Czech Digital Library. We selected the pages of the test subset so that the relative frequency of each category was similar to that of the entire set. To achieve this, we made several proposals of the test subset and selected one with the most similar distribution. The resulting test set contains a total of 1 93819381\,9381 938 annotated elements. The distribution of the test subset is shown in Figure 4(b).
The dataset is publicly available on Zenodo platform [^3] https://doi.org/10.5281/zenodo.12788419. It consists of pages from the Czech Digital Library, a text file containing a list of pages from the other datasets, and ground-truth text files in YOLO format for all pages. Since the test subset contains only pages from the Czech Digital Library, it is possible to use the dataset even without the need to obtain the other datasets.
@@ 4 Detection baselines
In this section, we present baseline methods and their results on the AnnoPage Dataset. We trained YOLO [13] and DETR [5] models on the development set and evaluated the best-performing models on the test set. More specifically, we split the development set into training and validation in the same manner as in the case of the test set, i.e. we aimed to obtain a similar distribution of the labels in the validation set as in the entire development set. As a result, we obtained a validation set of 509 pages with 1 92619261\,9261 926 annotated elements. First, we experimented with DETR and various YOLO models with fixed resolution of input images. We also investigated the effect of different image resolutions on the detector performance.
@@
The results of the trained models on the test set are summarized in Table 1. The results show that YOLO11m and YOLO11l perform the best and their results are very similar. However, the values of both metrics suggest that there is room for improvement. On the other hand, the DETR performed the worst, probably due to the fact that Transformers generally require very large datasets to optimize well [21, 23, 9].
In Figure 6, we present a confusion matrix for the YOLO11l model on the test set. The confusion matrix shows that the model most often confuses photographs (4) and cartoons (8) with images (12) and other technical drawings (14) with schemas (21). However, it should be mentioned that especially the cartoon is relatively less frequent in the test set, so any confusion is then even more pronounced in the matrix. The matrix also shows that the model quite often misses elements (B), especially chemical formulas (1), exlibris (3) and handwritten inscriptions (20).
Table 1: Results of object detection baselines. We report the mAP@50 and mAP@50-95 metrics of each model on the test set.
  Model mAP@50 mAP@50-95
  YOLO11n 0.616 0.558
  YOLO11s 0.641 0.585
  YOLO11m 0.658 0.598
  YOLO11l 0.657 0.602
  DETR 0.458 0.377
Figure 6: Confusion matrix for trained YOLO11m model. The numbers on the left and bottom indicate the categories defined in the dataset description (Section 3.2). The last row and column (labeled ’B’) is the category called ’Background’ – the bottom row represents false negatives and the last column represents false positives. The entire matrix is column-wise normalized.

### v2

@@ Abstract
We introduce the AnnoPage Dataset, a novel collection of 7,550 pages from historical documents, primarily in Czech and German, spanning from 1485 to the present, focusing on the late 19th and early 20th centuries. The dataset is designed to support research in document layout analysis and object detection. Each page is annotated with axis-aligned bounding boxes (AABB) representing elements of 25 categories of non-textual elements, such as images, maps, decorative elements, or charts, following the Czech Methodology of image document processing. The annotations were created by expert librarians to ensure accuracy and consistency. The dataset also incorporates pages from multiple, mainly historical, document datasets to enhance variability and maintain continuity. The dataset is divided into development and test subsets, with the test set carefully selected to maintain the category distribution. We provide baseline results using YOLO and DETR object detectors, offering a reference point for future research. The AnnoPage Dataset is publicly available on Zenodo [^1] https://doi.org/10.5281/zenodo.12788419, along with ground-truth annotations in YOLO format.
@@ 1 Introduction
In recent years, the digitization of documents in various institutions has made the content of documents accessible to the general public. However, while the textual content of documents is usually easily recognizable using today’s text recognition technologies (OCR), other document content is often not processed in libraries and archives, even though it may contain important information. Such content includes images, photographs, charts, various diagrams and drawings, musical notation, etc. The first and key step in processing such elements is the ability to detect them on the page and determine their category. Subsequently, these elements can be processed according to their category; for instance, images can be indexed for search, musical notations interpreted using optical music recognition (OMR), and mathematical expressions analyzed through math recognition systems. Such category-specific processing greatly expands the possibilities for users interacting with digitised documents, benefiting both general audiences and specialists conducting targeted research.
For document layout analysis and object detection in documents, many datasets exist [23, 33, 1, 30, 12, 9, 8, 18, 27, 6]. However, these datasets contain only several categories of elements [33, 21, 8], they focus mainly on identification of different parts of textual content, such as captions, titles, headers, footers, etc. [23, 19, 18], they are synthetically generated [21], or they focus only on modern documents [1, 33, 23, 19, 18]. Some datasets focus only on a single category, such as text lines [9, 15] or tables [25, 27, 17, 11]. Also, historical documents, preserved by numerous institutions, also contain unique visual elements typically absent from modern documents, such as friezes, vignettes, signets, and other decorative components.
In this paper, we present the novel AnnoPage Dataset, which contains 7 55075507\,5507 550 pages from various, mainly historical documents. The pages were written mostly in Czech and German at the end of the 19th and during the 20th century (see Figure 1 for example pages). These pages were carefully annotated by expert librarian annotators who used axis-aligned bounding boxes to identify elements of 25 categories of non-textual elements according to the Czech Methodology of image document processing [13].
Figure 1: Example pages from the AnnoPage Dataset with rendered ground-truth in colored bounding-boxes.
@@ 2 Related work
In general, document layout analysis is a task for which a large number of diverse datasets exist [15, 9, 28, 12, 23, 33, 6]. There are also datasets that focus on detecting various non-textual elements such as images, tables, or mathematical formulas [1, 8, 3, 7, 21, 4, 30, 33, 23, 19, 18]. In the terms of annotations, the ground-truth typically contains elements annotated either as bounding boxes [3, 4, 30, 33, 23, 17, 19] or bounding polygons [1, 8, 21, 7]. Some datasets contain pixel-level annotations, such as HBA 1.0 dataset [20], DIVA-HisDB [26], or U-DIADS-Bib [35].
One of the existing datasets designed for document layout analysis that has a different type of annotation is the , which contains pixel-level annotations.
@@ 3.1 Data acquisition
We have collected 7 55075507\,5507 550 pages from various documents. Most of the pages (5 69056905\,6905 690) come from publicly accessible documents from the Czech Digital Library [^2] https://www.digitalniknihovna.cz/. These pages are written mostly in Czech and German, but there are also a few pages containing other languages, such as French, and English. In terms of dating, the pages are from as early as 1485 to the present, with the majority from the late 19th and the first half of the 20th centuries. The distribution of the pages over the years is shown in Figure 2.
The rest of the pages come from existing datasets, most of them containing historical pages, namely IlluHisDoc [21] (237 pages), PRImA Layout Analysis Dataset [1] (170 pages), PRImA RDCL2019 [7] (40 pages), PRImA Europeana Newspapers [8] (361 pages), ICDAR2019 cBAD dataset [9] (382 pages), and TexBiG [30] (670 pages). The purpose of these extra pages is to increase the variability of the final dataset and also to maintain continuity across datasets. Although these datasets contain the original ground-truth, we have created new annotations for these pages following the Czech Methodology of image document processing [13]. To distinguish the Czech data from the pages from the extra datasets and the full dataset, we refer to the Czech data as the Czech subset in the rest of the paper.
@@ 3.2 Element categories
Figure 3: Examples of selected element categories, which are less common and mostly related to historical documents. These categories are highlighted by bounding-boxes in orange (Exlibris), green (Frieze), blue (Signet), and red (Vignette).
@@ 3.4 Final dataset
The final dataset contains 27 9042790427\,90427 904 annotated elements on 6 72667266\,7266 726 pages and 824 pages without any annotated element. Each element is represented by a tightly fitted axis-aligned bounding box (AABB). The distribution of the number of annotations for each category is shown in Figure 4(a). The dataset is divided into development and test subsets. The further division of the development subset into the traditional training and validation subsets is left to the users. The test subset contains 600 pages (21 pages without any annotated element) coming from the Czech subset. We selected the pages of the test subset so that the relative frequency of each category was as similar to the entire set as possible. The resulting test set contains a total of 1 93819381\,9381 938 annotated elements. The distribution of the test subset is shown in Figure 4(b).
The dataset is publicly available on Zenodo platform [^3] https://doi.org/10.5281/zenodo.12788419. It consists of pages from the Czech Digital Library, a text file containing a list of pages from the other datasets, and ground-truth text files in YOLO format for all pages. Since the test subset contains only pages from the Czech subset, it is possible to use the dataset even without the need to obtain pages from the extra datasets.
@@ 3.5 Limits of the dataset
During the annotation process, these challenging cases were discussed and decided among the expert librarian annotators with respect to the Czech methodology of image document processing.
@@ 4 Detection baselines
In this section, we present baseline methods and their results on the AnnoPage Dataset. We trained YOLO [14] and DETR [5] models on the development set and evaluated the best-performing models on the test set. More specifically, we split the development set into training and validation in the same manner as in the case of the test set, i.e. we aimed to obtain a similar distribution of the labels in the validation set as in the entire development set. As a result, we obtained a validation set of 509 pages with 1 92619261\,9261 926 annotated elements. To assess the contribution of the Czech data, we also trained models only on the Czech subset and compared them with models trained on the full dataset. First, we experimented with DETR and various YOLO models with fixed resolution of input images. We also investigated the effect of different image resolutions on the detector performance.
@@
The results of the trained models on the test set are summarized in Table 1. The results show that YOLO11m and YOLO11l perform the best and their results are very similar on both the Czech subset and the full dataset. When comparing the results between the Czech subset and the full dataset, it can be seen that the pages from the extra datasets bring improvements and help the models to generalize better by exposing them to a broader range of variability, although the Czech subset alone already provides solid baseline performance. However, the values of both metrics suggest that there is room for improvement. On the other hand, the DETR performed the worst, probably due to the fact that Transformers generally require very large datasets to optimize well [22, 24, 10].
Table 1: Results of object detection baselines. We report the mAP@50 and mAP@50-95 metrics on the test set of models trained on the Czech subset and the full dataset.
  Model Czech subset Full dataset
mAP@50 mAP@50-95 mAP@50 mAP@50-95
  YOLO11n 0.595 0.531 0.616 0.558
  YOLO11s 0.567 0.513 0.641 0.585
  YOLO11m 0.592 0.541 0.658 0.598
  YOLO11l 0.607 0.547 0.657 0.602
  DETR 0.438 0.356 0.458 0.377
In Figure 6, we present a confusion matrices for the YOLO11l model trained on the Czech subset and the full dataset and evaluated on the test set. The confusion matrix shows that the model most often confuses photographs (4) and cartoons (8) with images (12) and other technical drawings (14) with schemas (21). However, it should be mentioned that especially the cartoon is relatively less frequent in the test set, so any confusion is then even more pronounced in the matrix. The matrix also shows that the model quite often misses elements (B), especially chemical formulas (1), exlibris (3) and handwritten inscriptions (20).
(a) Czech subset
(b) Full dataset
Figure 6: Confusion matrices for YOLO11l models trained on the Czech subset (Fig. 6(a)) and the full dataset (Fig. 6(b)). The numbers on the left and bottom indicate the categories defined in the dataset description (Section 3.2). The last row and column (labeled ’B’) is the category called ’Background’ – the bottom row represents false negatives and the last column represents false positives. Each matrix is column-wise normalized.
Figure 7 contains example of pages with erroneous predictions of YOLO11l model trained on the full dataset. The examples on the left and in the middle show objects for which the categories were wrongly predicted. The sample on the right contains correctly identified categories (tables), but the model produced more fine grained detections than ground-truth. This is related to the limits of the dataset described in Section 3.5, where it is not clear whether it is a single table or multiple smaller tables.
Figure 7: Example of erroneous predictions produced by YOLO11l model trained on the full dataset. The first row contains pages with rendered predictions and the second row contains rendered ground-truth. Examples on the left and in the middle contain objects with wrongly identified classes while the right example contains more fine grained predictions than ground-truth.